The ratio if the measures of two complementary angles is 3:7. What is the measure of the larger angle?
Answer by jim_thompson5910(35256)

3x+7x = 90

10x = 90

x = 90/10

x = 9

3x = 3*9 = 27

7x = 7*9 = 63

The larger angle is 63 degrees.